The heat generation rate in a plane wall of 0.24 m thickness is 0.4 MW/m3

. The wall is exposed
on both sides to convection at 30°C. (a) Determine and compare the maximum temperatures
for k = 25, k = 50, k = 200 and k = 410 W/mK assuming h = 250 W/m2K. (b) Determine and
compare the maximum temperature for h = 50, 250, 500 and 1000 W/m2K with k = 25 W/mK.
Answer #1

(a) x Velunne h(2A) (Tw-30) arg qgx I convection connection 0.4x10ʻX AX (0.24)=2 ha ( TW-30) Tw 0.4/10*x A x0.24 = 2X250X Alfor max. temp. O -ergen + Cizo -org N ag = 0 ry n 24 10.24) 2 495+ 0:4806 dI da K (0:24) + k 2K K 2K = TE gan? rga tint C₂ f6 TE 495 +0.4810 [. 0.0144] 2x25 IT = 6.10.2k for k= 50 W/ bik T=495 + 0.LiX10C (0.0144] = 55206. 2 x 50 for k=200 w/mk T2495for h= 500 W/m²K Tw2:48.000 30 = 12.6°C 500 for h 21000 W/m²K Tw=48000 +30 = 78°c 100 T = 6 Maxinean temp here also at n=0, 1